OpenStudy (anonymous):

The OfficeJet printer can copy Janet's dissertation in 18 minutes. The LaserJet printer can copy the same document in 10 minutes. If the two machines work together, how long would they take to copy the dissertation?

OpenStudy (turingtest):

18x=1--->x=1/18 10y=1--->y=1/10 (x+y)t=1--->t=1/(x+y)=1/(28/180)=45/7=6 3/7 min

OpenStudy (turingtest):

or about 6.43 min